# BA-ORABE: Blockchain-Based Auditable Registered Attribute-Based Encryption With Reliable Outsourced Decryption

### v1

@@ 5.1 Static Security
Suppose the construction 4.3 of [8] is statically secure, then our sRABE with outsource decryption scheme 4.1 is also statically secure under KEM setting.
@@
Query. The simulator $\mathcal{B}$ answers the adversary $\mathcal{A}$’s key-generation queries on a non-corrupted slot index $i\in N$ with a set of attributes S. follows:
@@
Challenge. $\mathcal{A}$ submits a encapsulated key pair $(K_{0}^{*},K_{1}^{*})\in\{0,1\}^{2\times k}$ to $\mathcal{B}$, and $\mathcal{B}$ sends them to the challenger $\mathcal{C}$. The challenger $\mathcal{C}$ generate the challenge ciphertext pair $(CT_{0}^{*},CT_{1}^{*})$ corresponds to $(K_{0}^{*},K_{1}^{*})$ and sends it to $\mathcal{B}$.
@@
Suppose our sRABE with outsource decryption scheme 4.1 is statically secure, then our RABE with outsource decryption scheme 4.2 is also statically secure.
@@ 5.2 Verifiability
Assume that $H_{0}$, $H_{2}$ are collision-resistant hash functions, the RABE scheme with outsourced decryption is verifiable.
@@
Assume that there exists an adversary $\mathcal{A}$ that can break the verifiability of outsourced decryption, then we construct a simulator $\mathcal{S}$ can break the collision-resistance of hash functions $H_{0},H_{2}$. Given two challenging hash functions $H_{0}^{*},H_{2}^{*}$. $\mathcal{S}$ interacts with $\mathcal{A}$ as follows:
Setup: $\mathcal{S}$ runs Setup gives crs to $\mathcal{A}$.
Challenge $\mathcal{A}$ runs keygen generate $(sk^{*},pk^{*})$ and register (S*,pk*) to $\mathcal{S}$, $\mathcal{S}$ runs RegPK and gives mpk to $\mathcal{A}$.
Outputs $\mathcal{A}$ outputs an access structure $(M^{*},\rho^{*})$ and a tuple $\{CT^{*},sk^{*},CT_{1}^{*^{\prime}},CT_{2}^{*^{\prime}}\}$, where S* satisfies $(M^{*},\rho^{*})$. $\mathcal{A}$ wins the game if:
$Decrypt_{out}(sk^{*},CT_{1}^{*^{\prime}},CT^{*})\neq\perp\cap$
$Decrypt_{out}(sk^{*},CT_{2}^{*^{\prime}},CT^{*})\neq\perp\cap$
$Decrypt_{out}(sk^{*},CT_{1}^{*^{\prime}},CT^{*})\neq Decrypt_{out}(sk^{*},CT_{2}^{*^{\prime}},CT^{*})$
And then $\mathcal{S}$ can get:
@@
Table 1: Comparison of Pairing-Based RABE Schemes
@@
5.3 Exemptibility
Theorem 5.7.
Suppose that the soundness of $\prod_{DLEQ-NIZK}$ protocol holds, the RABE scheme with outsourced decryption is exempt.
Proof 5.8.
Assume that there exists an adversary $\mathcal{A}$ that can break the exemptibility of outsourced decryption, then we construct a simulator $\mathcal{S}$ can break the soundness of $\prod_{DLEQ-NIZK}$ protocol. Given challenging protocol $\prod_{DLEQ-NIZK}^{*}$. $\mathcal{S}$ interacts with $\mathcal{A}$ as follows:
Setup: $\mathcal{S}$ runs Setup gives crs to $\mathcal{A}$.
Challenge $\mathcal{A}$ runs keygen generate $(sk^{*},pk^{*})$ and register (S*,pk*) to $\mathcal{S}$, $\mathcal{S}$ runs RegPK and gives mpk and aux to $\mathcal{A}$.
Outputs $\mathcal{A}$ outputs an access structure $(M^{*},\rho^{*})$ and a tuple $\{CT^{*},CT^{*^{\prime}},hsk*,(sk_{1}^{*},\pi_{1}),(sk_{2}^{*},\pi_{2}),T_{i}^{*}\}$, where S* satisfies $(M^{*},\rho^{*})$. $\mathcal{A}$ wins the game if:
$hsk^{*}\leftarrow Update(crs,aux,pk^{*})\cap$
$CT^{*^{\prime}}\leftarrow Transform(hsk^{*},CT^{*})\cap$
$1\leftarrow\prod_{DLEQ-NIZK}^{*}.\textbf{Verify}(\pi_{1}^{\prime},(C_{2}^{\prime},C_{2}^{\prime sk_{1}^{*}},g,T_{i}^{*}))\cap$ $1\leftarrow\prod_{DLEQ-NIZK}^{*}.\textbf{Verify}(\pi_{2}^{\prime},(C_{2}^{\prime},C_{2}^{\prime sk_{2}^{*}},g,T_{i}^{*}))\cap$ $Decrypt_{out}(sk_{1}^{*},CT^{*^{\prime}},CT^{*})\neq Decrypt_{out}(sk_{2}^{*},CT^{*^{\prime}},CT^{*}).$ And then $\mathcal{S}$ can get
$sk_{1}^{*}\neq sk_{2}^{*}\ \cap$
$(C_{2}^{\prime},C_{2}^{\prime sk_{1}^{*}},g,T_{i}^{*})\in R_{DLEQ}\ \cap$
$(C_{2}^{\prime},C_{2}^{\prime sk_{2}^{*}},g,T_{i}^{*})\in R_{DLEQ}$
Thus $\mathcal{S}$ can break the soundness of $\prod_{DLEQ-NIZK}^{*}$ protocol.
5.4 Fairness
The fairness of our scheme ensures that the DCS gets paid if and only if it returns a correct result. The DU must deposit tokens into the smart contract to submit the outsourced decryption task. If the DCS submits the correct result, it will receive the reward at the end of the challenge window, which is ensured by exemptibility. Otherwise, DU will send valid fraud proof and get the deposit back, and DCS get nothing but pay for the transaction gas fee for malicious behavior. Thus the proposed scheme achieves the fairness property as long as the exemptibility and Ethereum consensus protocol is guaranteed.
Table 2: Comparison With Outsourced Decryption ABE Schemes
Schemes confidentiality verifiability exemptibility fairness decentralized outsourcing
[11]
[12]
[13]
[14]
[19]
[17]
Our Scheme
@@ 6.1 Theoretical Analysis
Table 1 shows the comparison of some existing Registered ABE schemes and our scheme. Since our scheme is based on [8], the parameters, except for the decryption cost, are identical to those of [8]. And the parameter sizes are better than other schemes. Besides, our outsourced decryption significantly reduces the final decryption cost to a constant level, requiring only one group exponentiation operation and one group multiplication operation.
Table 2 shows the functional comparison of the previous OABE schemes and our scheme in terms of confidentiality, verifiability, exemptibility, fairness, and decentralized outsourcing. Almost all schemes achieve confidentiality and verifiability. Only our scheme and [17] achieves exemptibility, fairness, and decentralized outsourcing property.
@@ 6.2 Experimental Evaluation
As we use the slotted registered ABE to registered ABE transformation[4], which constructs registered ABE through multiple slotted registered ABE instances, We only compared the basic slotted scheme here. We conducted experiment and performance analysis on the computation cost comparison between sRABE scheme in [8] and our sRABE with outsourced decryption scheme. We implemented them in Python language within the Charm framework. Our benchmark uses SS512 and SS1024 elliptic curves, SHA-256 hash function, and is executed on an Intel(R) Core(TM) i9-12900H CPU and 4 GB RAM running Ubuntu 20.04 LTS 64-bit and Python 3.8.10. For each algorithm, we repeat 1,000 times to get the average execution time.
As shown in Fig 4(a), the decryption time of sRABE scheme in [8] increases linearly with the numbers of attributes while our scheme with outsource decryption has constant decryption time and is 108$\times$-1585$\times$ in SS512 curve and 175$\times$-2647$\times$ in SS1024 curve lower than [8] in attribute numbers from 5 to 100, which is shown in Fig 4(b). The original and transform ciphertext size of our scheme are shown in Fig 4(c) and Fig 4(d) respectively. Note that we only consider RABE ciphertext and the symmetric encryption ciphertext is not included. The original ciphertext linearly increases with the numbers of attributes like [8] and the transform ciphertext contains only two $\mathbb{G}_{T}$ elements, which is about 0.25kb in SS512 curve and 0.5kb in SS1024 curve.
@@ 6.2.2 Fraud Proof Generation and Verification
We conducted the experiment and performance analysis on fraud proof generation and verification. The experiment setup and machine are the same as Section 6.2.1. The proving time and verifying time of fraud proof are about 0.87ms and 1.49ms in SS512 curve and 11.81ms and 21.40ms in SS1024 curve respectively as shown in Fig 5. It is important to note that fraud proofs are only required to be generated and verified during the dispute phase. Therefore, it does not constitute bottleneck for the system under our optimistic assumption.
@@
(a) publishState
(b) publishTag
(c) createTask
(d) submitResult
(e) publishFraudProof
(f) publishVerifyResult
Figure 6: Gas usage of (a) publishState, (b) publishTag, (c) createTask, (d) submitResult, (e) publishFraudProof and (f) publishVerifyResult for different numbers of attributes.
@@ 6.2.3 Gas Usage
7 Conclusions
In this paper, we proposed BA-ORABE, a fully auditable registered attribute-based encryption scheme with reliable outsourced decryption based on blockchain, and then presented a formal security proof of its static security, verifiability, and exemptibility. Our proposed scheme achieves verifiability through a verifiable tag mechanism and exemptibility through NIZK proof. Furthermore, We extended fully auditable property and realized fairness and decentralized outsourcing based on public blockchain. Finally, we compared the theoretical complexity of BA-ORABE with current existing schemes and implemented and evaluated it on Ethereum to demonstrate efficiency and feasibility.

### v2

@@ 5.1 Static Security
Suppose the construction 4.3 of [8] is statically secure, then our sRABE with outsourced decryption scheme 4.1 is also statically secure under KEM setting.
@@
Query. The simulator $\mathcal{B}$ answers the adversary $\mathcal{A}$’s key-generation queries on a non-corrupted slot index $i\in N$ with a set of attributes S as follows:
@@
Challenge. $\mathcal{A}$ submits a plaintext pair $(K_{0}^{*},K_{1}^{*})\in\{0,1\}^{2\times k}$ to $\mathcal{B}$, and $\mathcal{B}$ sends them to the challenger $\mathcal{C}$. The challenger $\mathcal{C}$ generate the challenge ciphertext pair $(CT_{0}^{*},CT_{1}^{*})$ corresponds to $(K_{0}^{*},K_{1}^{*})$ and sends it to $\mathcal{B}$.
@@
Suppose our slotted registered ABE with outsourced decryption scheme 4.1 is statically secure, then our registered ABE with outsourced decryption scheme 4.2 is also statically secure.
@@ 5.2 Verifiability
Assume that $H_{0}^{*}$, $H_{2}^{*}$ are collision-resistant hash functions, the registered ABE scheme with outsourced decryption is verifiable.
@@
The verifiability is broken if DCS can return two different transform ciphertexts to match original tag. Then $\mathcal{S}$ can get:
@@
5.3 Exemptibility
Theorem 5.7.
Suppose that the soundness of $\prod_{DLEQ-NIZK}$ protocol holds, the registered ABE scheme with outsourced decryption is exempt.
Proof 5.8.
The exemptibility is broken if DU can return two different valid verification keys $vk_{1}$ and $vk_{2}$ in fraud proof. Then $\mathcal{S}$ can get:
$1\leftarrow\prod_{DLEQ-NIZK}.\textbf{Verify}(\pi_{1},(C_{2}^{\prime},vk_{1},g,T_{i}))$
$1\leftarrow\prod_{DLEQ-NIZK}.\textbf{Verify}(\pi_{2},(C_{2}^{\prime},vk_{2},g,T_{i}))$
where
$vk_{1}\neq vk_{2}.$
Thus $\mathcal{S}$ can break the soundness of $\prod_{DLEQ-NIZK}$ protocol.
5.4 Fairness
Theorem 5.9.
The registered ABE scheme with outsourced decryption based on Ethereum keeps fairness.
Proof 5.10.
The fairness of our scheme ensures that the DCS gets paid if and only if it returns a correct result. The DU must deposit tokens into the smart contract to submit the outsourced decryption task. If the DCS submits the correct result, it will receive the reward at the end of the challenge window, which is ensured by exemptibility. Otherwise, the malicious behavior will be detected by DU, which is ensured by verifiability, and DCS get nothing but pay for the transaction gas fee for malicious behavior. Thus the proposed scheme achieves the fairness property as long as the verifiability, the exemptibility, and Ethereum consensus protocol is guaranteed.
Table 1: Comparison of Pairing-Based Registered ABE Schemes
@@ 6.1 Theoretical Analysis
Table 1 shows the comparison of some existing registered ABE schemes and our scheme. Since our scheme is based on [8], the parameters, except for the decryption cost, are identical to those of [8]. And the parameter sizes are better than other schemes. Besides, our outsourced decryption significantly reduces the final decryption cost to a constant level without any pairing operation, requiring only one group exponentiation operation.
Table 2 shows the comparison of the previous OABE schemes and our scheme in terms of authority, verifiability, exemptibility, fairness, decentralized outsourcing, and final decryption cost. Our scheme is the only one that does not rely on a trusted authority and achieves the lowest decryption overhead among all the compared schemes. While most schemes ensure verifiability, only our scheme and [17] provide exemptibility, fairness, and decentralized outsourcing capabilities.
@@ 6.2 Experimental Evaluation
As we use the slotted registered ABE to registered ABE transformation[4], which constructs registered ABE through multiple slotted registered ABE instances, We only compared the basic slotted scheme here. We conducted experiment and performance analysis on the computation cost comparison between sRABE scheme in [8] and our sRABE with outsourced decryption scheme, excluding the computational cost of symmetric encryption for consistency. We implemented them in Python language within the Charm framework. Our benchmark uses SS512 and SS1024 elliptic curves, SHA-256 hash function, and is executed on an Intel(R) Core(TM) i9-12900H CPU and 4 GB RAM running Ubuntu 20.04 LTS 64-bit and Python 3.8.10. For each algorithm, we repeat 1,000 times to get the average execution time.
@@
As shown in Fig 4(a), the decryption time of sRABE scheme in [8] increases linearly with the numbers of attributes while our scheme with outsourced decryption has constant decryption time and is 108$\times$-1585$\times$ in SS512 curve and 175$\times$-2647$\times$ in SS1024 curve lower than [8] in attribute numbers from 5 to 100, which is shown in Fig 4(b). The original and transform ciphertext size of our scheme are shown in Fig 4(c) and Fig 4(d) respectively. Note that we only consider registered ABE ciphertext and the symmetric encryption ciphertext is not included. The original ciphertext linearly increases with the numbers of attributes like [8] and the transform ciphertext contains only two $\mathbb{G}_{T}$ elements, which is about 0.25kb in SS512 curve and 0.5kb in SS1024 curve.
Table 2: Comparison With Outsourced Decryption ABE Schemes
Schemes Authority-Free Verifiability Exemptibility Fairness Decentralized Outsourcing Decryption Cost
[…]
[12] $\mathbb{E}_{T}$
[13] $\mathbb{E}_{T}$
[…]
[17] $\mathbb{E}_{T}$
Our Scheme $\mathbb{E}_{T}$
@@ 6.2.2 Fraud Proof Generation and Verification
We conducted the experiment on fraud proof generation and verification. The experiment setup and machine are the same as Section 6.2.1. The proving time and verifying time of fraud proof are about 0.87ms and 1.49ms in SS512 curve and 11.81ms and 21.40ms in SS1024 curve respectively as shown in Fig 5. It is important to note that fraud proofs are only required to be generated and verified during the dispute phase. Therefore, it does not constitute bottleneck for the system under our optimistic assumption.
@@ 6.2.3 Gas Usage
Figure 6: Gas usage of different contract functions
7 Application
Decentralized Finance(DeFi) [39] signifies a fundamental shift from traditional banking and financial services, providing a blockchain-based alternative that eliminates the reliance on central authorities. By utilizing smart contracts on the Ethereum blockchain, DeFi allows users to lend, borrow, trade assets, and earn interest within a permissionless and transparent ecosystem. Applying ABE in DeFi can provide users with fine-grained access control for transactions, enhancing the security of digital asset trading in decentralized exchanges. Ma et al.[40] proposed a scheme based on traditional ABE where users define attributes such as asset type, amount, exchange rate, and time, and set diverse access control policies. Only transaction counterparts that satisfy the user’s specified policies are allowed to engage in the trade, thereby enhancing transaction security and ensuring compliance with user preferences in decentralized exchanges.
However, traditional ABE schemes rely on one or multiple trusted authorities, which still poses certain security concern when applied to DeFi. In contrast, registered ABE eliminates the need for any trusted authority, and its transparent process makes it more suitable for application in DeFi. Our auditable scheme further enhances the security of the system. Moreover, due to the widespread use of mobile devices for transactions, the high decryption overhead in ABE presents a significant challenge. The outsourced decryption registered ABE scheme effectively addresses this issue. In summary, the auditable registered ABE scheme with outsourced decryption proposed in this work provides significant advantages for transaction control in DeFi.
8 Conclusion
In this paper, we proposed BA-ORABE, a fully auditable registered attribute-based encryption scheme with reliable outsourced decryption based on blockchain, and then presented security proof of its static security, verifiability, and exemptibility. Our proposed scheme achieves verifiability through verifiable tag mechanism and exemptibility through NIZK proof. Furthermore, We extended fully auditable property and realized fairness and decentralized outsourcing based on public blockchain. Finally, we compared the theoretical complexity of BA-ORABE with current existing schemes and implemented and evaluated it on Ethereum to demonstrate efficiency and feasibility. We also showed that our scheme provides significant advantages in DeFi application, which could be explored further in future work.